Note. Apply Discrete Probability Distributions to solve the given problems

Q1. The experience of a house agent indicates that he can provide suitable accomodation for 75% of the clients who come to him. If on a particular occasion, 6 clients approach him indepndently, calculate the probability that   

  1. Less than 4 clients
  2. Exactly 4 clients
  3. At least 5 clients, will get satisfactory accommodation 
1
Expert's answer
2021-08-27T10:26:09-0400

Suppose we let X = the number of clients approaching the house agent who get satisfactory accommodation, then the random variable X follows a binomial distribution with n=6 and p=0.75

Given the number of trials, n, and the probability of success, p,

P(X=x) = "\begin{pmatrix}
 n \\
 X 
\end{pmatrix}"px(1-p)n-x for x = 0, 1,2,3,4,5 and 6


1.Less than 4 clients


Given n=6 and p=0.75

P(X<4) = P(x=0) +P(x=1) + P(x=2) +P(x=3)

P(x=0) = "\begin{pmatrix}
 6 \\
 0 
\end{pmatrix}"0.750(1-0.75)6-0 = 0.0002 

P(x=1) = "\begin{pmatrix}
 6 \\
 1 
\end{pmatrix}"0.751(1-0.75)6-1 = 0.0044

P(x=2) =   "\begin{pmatrix}
 6 \\
 2 
\end{pmatrix}"0.752(1-0.75)6-2 = 0.0330

P(x=3) = "\begin{pmatrix}
 6 \\
 3 
\end{pmatrix}"0.753(1-0.75)6-3 = 0.1318

Thus,

P(X<4) = 0.0002 + 0.0044 +0.0330 + 0.1318 = 0.1694

Therefore, the probability that less than 4 clients will get satisfactory accommodation is approximately equal to 0.1694


2.Exactly 4 clients


P(4) =   "\begin{pmatrix}
 6 \\
 4 
\end{pmatrix}"0.754(1-0.75)6-4 = 0.2966

The probability that exactly four clients will get satisfactory accommodation is approximately equal to 0.2966


3.At least 5 clients, will get satisfactory accommodation


P(X≥5) = P(5) + P(6)

P(5) = "\begin{pmatrix}
 6 \\
 5 
\end{pmatrix}"0.755(1-0.75)6-5 = 0.3560

P(6) =   "\begin{pmatrix}
 6 \\
 6 
\end{pmatrix}"0.756(1-0.75)6-6 = 0.1780

Thus,

P(X≥5) = P(5) + P(6) = 0.3560 + 0.1780 ≃ 0.5340

Therefore, the probability that at least 5 clients will get satisfactory accommodation is approximately equal to 0.5340